OpenStudy (wiggler):

limit as x tends to infinity of ln(x^2 + 2)/sqrt(x)

OpenStudy (wiggler):

\[\lim_{x \rightarrow \infty} \frac{ x^2+2 }{ \sqrt{x} }\]Show working.

OpenStudy (anonymous):

\[\lim_{x\to\infty}\frac{\ln(x^2+2)}{\sqrt x}=\frac{\infty}{\infty}\] Apply L'Hopital's rule: \[\lim_{x\to\infty}\frac{\frac{2x}{x^2+2}}{\frac{1}{2\sqrt x}}=\lim_{x\to\infty}\frac{4x^{3/2}}{x^2+2}=\frac{\infty}{\infty}\] Apply again: \[\lim_{x\to\infty}\frac{6x^{1/2}}{2x}=\frac{\infty}{\infty}\] One more time: \[\lim_{x\to\infty}\frac{3}{2\sqrt x}\]

OpenStudy (wiggler):

So as x-> infinity,\[\frac{ 3 }{ 2\sqrt{x} }\rightarrow0\]

OpenStudy (wiggler):

Is that right?

OpenStudy (anonymous):

Yes

OpenStudy (wiggler):

Thank you!

OpenStudy (anonymous):

yw